OpenStudy (anonymous):

What is the range of the graph of \[y = –3(x – 4)^2 + 1\] \[y\le 3\] \[y\ge 3\] \[y \le 1\] \[y\ge 1\]

OpenStudy (anonymous):

@Hero

hero (hero):

Hint: Expand \((x-4)^2\)

OpenStudy (anonymous):

(x-4)(x-4)...?

OpenStudy (anonymous):

use the following for expansion \[\Large (a-b)^=a^2+b^2-2ab\]

hero (hero):

You forgot the squared symbol

OpenStudy (anonymous):

\[x^2-8x+16\]

hero (hero):

Now multiply that by 3 and add 1

OpenStudy (anonymous):

\[-3x^2+24x-49\]

hero (hero):

Now set \(y = -3x^2 + 24x - 49\) then use the vertex formula \(x = -\large\frac{b}{2a}\) to find the x value

hero (hero):

After you find the x value, find the y value.

OpenStudy (anonymous):

\[y=-1\]

hero (hero):

Are you 100% sure about that?

hero (hero):

I think you missed a sign when calculating

OpenStudy (anonymous):

how do i know if its \[y\ge~or~~y\le\]

OpenStudy (anonymous):

@Hero why not to make life easier and just put x=4 you will be left with 1 only ?

hero (hero):

True

OpenStudy (anonymous):

okay but how do i know which sign?? :\

hero (hero):

Funny thing is, I keep getting y = -1

OpenStudy (anonymous):

yes exactly O.O

OpenStudy (anonymous):

yes that's good question. multiply the whole equation with minus one (your orignal equation) always try to make the square term positive always look for the sign of variable which is linear (in this case y) so when you multiply with -1 the y will be negative it means parabola will open towards -y axis

OpenStudy (anonymous):

which means....??

OpenStudy (anonymous):

put x=4 everything will be zero you will be left with 1 only. this means \[\Large y \le1\]

hero (hero):

I see. He means to plug x = 4 back into the original equation to get y = 1

hero (hero):

That let's you know that the y coordinate of the vertex is 1. The graph is concave down so the y values will go more negative and the graph progresses, which means that \(y\le1\)

OpenStudy (anonymous):

here is helpful graph. you can see here y will never be greater than 1 also parabola is open downward

hero (hero):

Also, I just realized that your equation should have been \(-3x^2 + 24x - 47\) You added 1 rather than subtracted 1. That's why we kept getting y = -1

OpenStudy (anonymous):

oh sorry